If Joan read 75 pages in 4 hours how long will it take her to read 250 pages?
Answer by stanbon(75887)

If Joan read 75 pages in 4 hours how long will it take her to read 250 pages?
------
x/250 = 4/75
---
x = 250(4/75)
---
x = 1000/75
x = 13 1/3 hrs
-------------------
Cheers,
Stan H.
========================